• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name

    Log entry time 16:22:00 on March 22, 2014

    Entry number 387247

    keyword=Status

    Day Shift starting
    Workers
    
    Day Shift Summary
    Workers
    Lindgren (SL)
    Kalyan RC  (not here much)
    Keppel TO
    Hashir 3rd
    
    Shift started waiting for pulsed beam.
    10:00 AM MCC request turn Moller Magnets off
    Only have optics tune with Moller magnets off.
    Kalyan and Thia enter Hall to turn Moller Magnets off
    Lost NMR Lock on LHRS
    Jack came in to diagnose problem.
    Problem is on IOC on Accelerator software side.
    Probably have to go into Hall to reset control voltage.
    Mike was called in by MCC crew to fix IOC problem
    We can still take beam without NMR lock.
    1:00 PM A few minutes of pulsed beam delivered.
    Suddenly lost beam.
    MCC called to report a gate valve problem  - lost control of valve
    1:30 PM Called more people in to fix it.
    Mike from MCC called in to report IOC fixed. 
    To check it we have to reset the voltage in the Hall
    MCC called still need  more time to fix valve.
    Decide to enter Hall to power cycle NMR
    3:45 Hall A controlled Access
    Thia and  Vince try to power cycle NMR meter
    Report back there is no power to the NMR meter.
    4:10 PM Swing shift taking over
    Vince will make full report
    
    Summary by shift leader